OpenStudy (anonymous):

38.5^x = 6.5^(x-2) how can i find x using logarithm? Help please.

OpenStudy (campbell_st):

take the log of both sides and use this log law \[\log(a^b) = b \times \log (a)\] then collect like terms and solve

OpenStudy (anonymous):

ok. i'll try.

OpenStudy (anonymous):

awesome, i got it, thanks a lot! :)